Circular No. 9008

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 Y2 (GIBBS) A. R. Gibbs reports his discovery of a comet with a narrow 16′′–20′′ tail

in p.a. 280^{o}–285^{o} and a well-condensed coma of size 13′′.5 × 9′′ on co-added unfiltered CCD images obtained with the Catalina Sky Survey (CSS) 0.68m Schmidt telescope (discovery observation tabulated below). E. Beshore (Mt. Lemmon 1.5-m reflector, 2009 Jan. 1.46–1.48 UT) adds that four coadded 30-s exposures show a condensed nuclear condensation with a tail ∼ 1′ long in p.a. ≈ 280^{o}. T. Spahr, Smithsonian Astrophysical Observatory, identified this comet with pre-discovery astrometry in the Minor Planet Center (MPC) archive from 2008 Dec. 1 (Mt. Lemmon) and 21 (CSS). Following posting on the MPC’s ‘NEOCP’ webpage, several other CCD astrometrists have noted the object’s cometary appearance. J. V. Scotti writes that Spacewatch images taken with the 1.8-m f/2.7 reflector show a 9′′ coma and a 0′.48 tail in p.a. 284^{o} on 2009 Jan. 1.34–1.35. R. Ligustri (Udine, Italy) co-added ten 180-s exposures taken remotely with a 0.25-m f/3 reflector near Mayhill, NM, U.S.A., on Jan. 1.44–1.45 to see a coma of diameter ∼ 15′′; twenty-four 120-s co-added images obtained similarly by E. Guido, G. Sostero, and P. Camilleri on Jan. 1.45–1.48 show a compact coma of diameter ∼ 10′′ and a narrow 10′′-long tail in p.a. 270^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 31.36701  10^{h}04^{m}12.^{s}21  +19^{o}49′55′′.3  18.2

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2009-A3.

T = 2009 Jan. 22.4628 TT  ω = 162.^{o}3785
e = 0.543639  Ω = 330.8890 }2000.0
q = 1.638444 AU  i = 7.2768
a = 3.590233 AU n^{o} = 0.1448839 P = 6.80 years

COMET P/2008 X4 (CHRISTENSEN) K. Kadota, Ageo, Japan, has reported astrometry for this comet (cf.

IAUC 9005) from CCD images obtained with a 0.25-m f/5 reflector on 2008 Dec. 14.35 (at low altitude in very bright twilight), 30.9, and 31.9 UT, for which he provided total magnitudes 11, 10.6, and 10.5, respectively. The astrometry for Dec. 30 and 31, together with improved orbital elements by B. G. Marsden and an ephemeris, appear on MPEC 2009-A1.

2009 January 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green